# Carolina Torreblanca

### v1

@@ Abstract
How has the credibility revolution reshaped political science? We address this question by using a large language model to classify 91,632 articles published between 2003 and 2023 across 174 political science journals, focusing on causal research designs, transparency practices, and citation patterns. Design-based studies—research strategies that explicitly a research design and the assumptions required for causal identification—have become increasingly common, displacing regression-based analyses that rely primarily on modeling assumptions. Yet as of 2023, studies without an explicit identification strategy still constitute nearly 40% of empirical quantitative work. Within design-based research, survey experiments dominate, while field experiments and quasi-experimental approaches have grown more modestly. Transparency practices such as placebo tests and power analysis remain rare. Design-based studies are concentrated in top journals and among authors at highly ranked institutions, and enjoy a persistent citation premium. The credibility revolution has meaningfully reshaped the discipline, though unevenly and incompletely.
@@ Introduction
Over the last two decades, political science has been swept by what many describe as the credibility revolution: an intellectual movement advocating a change in how researchers conceptualize and conduct causal empirical scholarship (Angrist and Pischke 2010; Dunning 2012; Gerber and Green 2012; Morgan and Winship 2014; Samii 2016; Aronow and Miller 2019; Blair, Coppock and Humphreys 2023; Slough and Tyson 2024; Imbens and Xu 2025). Or has it? In this paper, we examine whether and to what extent the credibility revolution has altered the landscape of empirical causal research in political science.
@@
As an epistemic movement that reshapes what researchers view as credible evidence when assessing empirical, especially causal, claims, the credibility revolution carries genuine transformative potential. Its intellectual core is a reconceptualization of causality. It encourages a shift away from relying primarily on model-based strategies that impose strong functional form assumptions and often implicit views of treatment assignment, toward design-based approaches that exploit how treatments are assigned and rest on explicit identification assumptions. This movement views credibility not solely as a consequence of adopting specific designs such as experiments or natural experiments; additionally, it proposes a clearer articulation of identification assumptions and supporting evidence through transparency practices such as pre-analysis plans, power analyses, and placebo tests (Sekhon 2009). At the same time, critics have raised concerns about external validity and over-reliance on as-if-random designs, underscoring that research design and substantive relevance must be considered jointly. Whether, on balance, this movement has produced an epistemic revolution in political science depends on how widely and deeply these ideas have reshaped actual research practice.
@@
Quantifying the depth and breadth of this potential transformation presents two measurement challenges. The first is scope. What constitutes a sufficiently representative sample of political science to allow statements about the discipline as a whole? Previous research examining methodological changes in other social science disciplines has focused on elite journals or publishing outlets (e.g., Currie, Kleven and Zwiers 2020; Goldsmith-Pinkham 2024). However, the discipline extends far beyond these outlets, which may publish systematically different research than the rest of the field. Consequently, systematic assessment requires analyzing thousands of papers across multiple journal tiers over a sufficient time period to capture potential changes. We address this by analyzing 91,632 articles from 174 political science journals spanning 2003–2023 using large language models (LLMs). This represents, to our knowledge, the most comprehensive analysis of political science research assembled to date.
@@
The second is operationalization. The credibility revolution is an intellectual movement that seeks to transform quantitative empirical research. Translating such ideas into concrete research decisions is left to individual researchers. What changes should we expect to observe if the revolution has taken hold? We focus on three dimensions where such transformation should be evident: the methodological approaches employed in empirical quantitative political science papers, the adoption of practices that enhance credibility beyond research design, and the professional rewards these approaches receive relative to the approaches they aim to substitute.
@@
We leverage advances in LLMs to classify our corpus along these dimensions. The strategy first maps papers by methodology and research goals, then focuses on explanatory empirical quantitative papers, the subset most likely to be shaped by the credibility revolution. Within this subset, we distinguish between design-based approaches that embody the principles of the credibility revolution and traditional model-based approaches, including linear regression and more complex statistical models. We also code the use of ancillary transparency practices, such as pre-analysis plans, power analysis, placebo tests, and robustness checks.
@@
We first assess the extent to which design-based empirical research has substituted model-based approaches in quantitative studies that aim to explain social phenomena, many of which make explicit causal claims. We find substantial—though uneven—substitution. Design-based methods increased from 15% to 40% of explanatory quantitative studies between 2003 and 2023, while model-based approaches declined from 57% to 39%. Yet this rise does not amount to a wholesale replacement. It is driven largely by survey experiments, which accounted for 44% of all design-based papers in 2023, and it remains concentrated in top journals. Overall, the methodological shift is real but circumscribed: traditional strategies—often relying on assumptions that remain unstated or untested—continue to play a central role in explanatory political science research.
@@
Second, we examine the adoption of practices that enhance credibility beyond research design. Such transparency practices remain uncommon even among design-based studies. Pre-analysis plans and power analyses appear in only 6% and 16% of experimental papers, respectively, and placebo tests are reported in just 22% of design-based observational studies. These patterns indicate that the credibility revolution’s influence beyond methodological substitution remains limited and has not yet generated broader shifts in research standards.
@@
Third, we assess whether design-based research receives different professional rewards than model-based approaches. Descriptively, we find evidence of a citation premium for design-based studies: such papers receive more citations than model-based papers, and this gap has widened over time. The premium has grown but now appears to have stabilized, suggesting that the discipline’s reward structure has shifted to favor more credible methods and may be reaching a new equilibrium.
@@
Overall, we find evidence of a partly successful revolution. Design-based methods have grown while model-based approaches have declined, indicating an increase in credible quantitative research. Yet the transformation is neither wide nor deep. It is not wide for two reasons. First, much of political science lies outside the domain where the credibility revolution’s lessons most directly apply: quantitative studies that seek to identify causal relationships. This is not inherently problematic, as high-quality descriptive research remains essential (De Kadt 2025). Second, traditional model-based approaches remain deeply embedded in researchers’ practices. It is not deep in the sense that, even among scholars adopting design-based methods, broader transparency practices remain rare, and the shift is driven largely by one approach–survey experiments, which accounted for nearly half of all design-based papers in 2023. Thus, the credibility revolution in political science is best understood as a reform: a real but limited transformation concentrated in relatively small segments of the discipline.
@@
Our paper builds on Currie, Kleven and Zwiers (2020) and Goldsmith-Pinkham (2024), who quantify the credibility revolution in economics using papers published in elite journals and National Bureau of Economic Research (NBER) working papers. A key limitation of these studies is that NBER affiliates do not represent the economics discipline as a whole. By contrast, we analyze articles from all political science journals with an impact factor above one, allowing us to examine the field well beyond elite institutions. We also extend Garg and Fetzer (2025), who use LLMs to identify causal claims in economics, by systematically classifying design choices and transparency practices across an entire discipline. Finally, our work contributes to a broader literature on how the credibility revolution has shaped political science, including patterns of knowledge production (Grossman, Dinneen and Torreblanca 2025), graduate training and hiring, and the evolution of methodological norms.
@@
We use the term “credibility revolution” to refer to the movement in the social sciences that sought to change how empirical causal research is conducted and evaluated. The motivating ideas marked a clear departure in how causality is conceptualized, empirically investigated, and defended.
The credibility revolution began in applied microeconomics in the early 1990s, driven by natural experiment studies that exploited institutional rules, policy changes, and quasi-random variation to identify causal effects (e.g., Card and Krueger 1994; Angrist and Krueger 1991). These empirical innovations spurred renewed emphasis on, and further development of, the potential outcomes framework and related causal theories.
Before this shift, much quantitative causal research in economics and political science relied on parametric regression models that treated causal and statistical relationships similarly. Researchers typically modeled the conditional expectation of an outcome using treatment variables and covariate (often not clearly separated) and then interpreted regression coefficients as causal effects. Identification assumptions were implicit, often resting on functional-form restrictions or conditional independence claims that were difficult to justify empirically (Angrist and Pischke 2010).
In what follows, we distinguish between two broad approaches to causal inference:
  • Design-based approaches identify causal effects by exploiting a clearly specified research design that provides (or approximates) exogenous variation in treatment assignment and by stating identification assumptions explicitly and separately from functional-form or distributional assumptions.
  • Model-based approaches identify causal effects primarily through parametric or statistical modeling assumptions—such as linearity, additivity, conditional independence, and distributional assumptions—often interpreting regression coefficients as causal effects without an explicit research design or separately articulated identification strategy.
One theoretical underpinning of the credibility revolution is the Neyman–Rubin potential outcomes framework (Rubin 1974; Holland 1986; Imbens and Rubin 2015). In this framework, causal effects are defined as contrasts between a unit’s potential outcomes under different treatment conditions, and causal inference is framed as a missing-data problem. A complementary development is the structural causal model and directed acyclic graph framework, which provides a graphical representation of causal assumptions and identification strategies (Pearl 2009). A central contribution from this literature is the local average treatment effect framework (LATE), which clarifies what instrumental variables identify and under what assumptions (Imbens and Angrist 1994).
Design-based approaches identify causal effects by exploiting randomization, institutional rules, policy cutoffs, or timing that generate exogenous variation and justify explicit identification assumptions (Imbens and Rubin 2015). Natural experiments, randomized controlled trials, instrumental variables, regression discontinuity designs, and difference-in-differences designs are canonical examples. In these designs, assumptions required for identification are stated explicitly, tied to empirical context, and conceptually separated from functional-form or distributional assumptions.
Two designs played especially influential roles. Regression discontinuity, first proposed by Thistlethwaite and Campbell (1960) and formalized for modern use by Lee and Lemieux (2010), became a central tool for credible analysis near institutional cutoffs. Difference-in-differences emerged as another pillar (Card and Krueger 1994). Although the parallel trends assumption can be seen as a functional form restriction, it can also be viewed as an ignorability condition applied to first differences or demeaned outcomes. We treat difference-in-differences as design-based because its use requires researchers to state the research design and identification assumptions clearly.
Political science followed a similar trajectory. Foundational work on field experiments (Gerber and Green 2000, 2012) and the synthesis of natural experiments in Dunning (2012) helped institutionalize design-based reasoning and advance credibility norms across the discipline. Because of the central role of electoral studies, regression discontinuity became a powerful and widely used tool (Eggers et al. 2015). LATE remains central to modern design-based instrumental-variable analysis in political science (Sovey and Green 2011; Lal et al. 2024). Difference-in-differences is also widely applied, though recent work highlights challenges posed by heterogeneous treatment effects, staggered adoption, and treatment reversals, all common in political science settings (Chiu et al. 2025). Together, these developments reshaped what counts as credible causal evidence in the field.
The emphasis on credibility extends beyond research design. Robustness and sensitivity analyses—placebo tests, balance checks, manipulation tests, and formal sensitivity procedures—assess how results respond to departures from identifying assumptions (Sekhon 2009; Blair, Coppock and Humphreys 2023; Cinelli and Hazlett 2020). At the same time, transparency practices such as pre-analysis plans, power analyses, open data and code, and clear reporting of deviations from pre-registered plans have become increasingly common. Journals and professional associations have institutionalized these norms through data and code policies and reproducibility checks, embedding transparency as a central component of credible empirical practice.
Debates accompany this shift. Critics argue that design-based estimands can be unduly local or may emphasize identification over substantive relevance (e.g., Heckman and Smith 1995) or external validity (Findley, Kikuta and Denly 2021). Others question strong assumptions behind randomized-control trials or as-if random designs (e.g., Deaton and Cartwright 2018). These debates shape how credibility is interpreted across fields and underscore that research design and theoretical relevance are complementary.
@@
The core ideas of the credibility revolution generate testable empirical predictions about how quantitative explanatory research should change. We focus on three dimensions: (a) methodological practices, (b) transparency and research standards, and (c) professional incentives.
First, if the credibility revolution has taken hold, design-based methods should increasingly displace purely model-based approaches in quantitative explanatory research, with growth spread across multiple design-based strategies.
Second, the credibility revolution should lead to greater use of transparency practices that strengthen the credibility of causal claims. Pre-analysis plans, power analyses, placebo tests, and transparent reporting of data and code should become more prevalent, as each addresses a distinct but complementary credibility concern.
Third, the credibility revolution implies a shift in the standards by which causal research is evaluated, which in turn should be reflected in disciplinary rewards. If the field increasingly views design-based approaches and transparency practices as markers of credible inference, then research employing these strategies should attract greater attention. As a result, design-based studies would be expected, on average, to receive more citations and to be published more frequently in prominent outlets than otherwise comparable model-based work.
@@
We explore these predictions using a corpus of political science research from the past two decades. In this section, we describe the data, key measurements, and methods used to construct these measures.
Data.
We begin with a set of 188 journals identified by Clarivate as part of the political science discipline and having a Scientific Journal Rankings (SJR) score of at least one. This threshold roughly corresponds to the upper tier of the field in citation influence and focuses attention on outlets that sit near the center of political science’s citation network. From this list, we removed journals for three distinct reasons: indexing (three journals not indexed in Scopus), quality control (three journals that lacked peer review), and language or format (seven journals that were not published in English and one book series). The final sample contains 174 English-language, peer-reviewed, Scopus-indexed political science journals. Our aim is not to cover every outlet, but to assemble a comprehensive corpus of articles that shape debates in the discipline.
@@
Using Scopus, we extracted journal-level metadata, including citation statistics, and article-level information for all publications in the selected journals. This process produced a dataset of 129,751 articles published between 2003 and 2023. To systematically code research design choices, we obtained full-text versions of the articles where possible and successfully retrieved the complete text for 91,632 articles from 156 journals. Full-text availability varies somewhat across publishers and subfields, so the full-text corpus is broad but not exhaustive. We then used a combination of supervised machine learning and a commercial LLM, gpt-4o, to classify articles along 19 research design dimensions (see description of the approach below). Figure 1 shows the yearly number of published papers in our full-text sample, both overall and by subfield. The volume of published work has grown rapidly. From 1,546 papers in 2003, the count has more than quadrupled to over 8,109 in 2023.
Key variables.
We construct a series of variables to characterize each paper’s methodological features, substantive focus, and transparency practices. Details of the coding procedure appear in the Supplementary Materials.
@@
We first assign each paper a subfield label drawn from six categories: American Politics, Comparative Politics, International Relations, Methodology and Formal Theory, Political Theory and Philosophy, and Public Policy/Administration. The label reflects the paper’s primary substantive contribution.
@@
We then identify whether the paper is an empirical quantitative study. A paper is coded as empirical quantitative if it conducts its own analysis of observational or experimental data, including reanalyses of existing datasets. Simulation exercises or purely methodological demonstrations without data are not included. Papers whose main contribution is qualitative, formal-theoretic, or normative are coded as non-quantitative for our purposes, though we distinguish these categories in the underlying coding. We assign general goals only to empirical quantitative papers, because only these studies contain measured relationships between variables that can be classified consistently as descriptive, explanatory, or predictive.
@@
Explanatory papers investigate the causes or consequences of social phenomena and develop evidence consistent with one or more causal relationships, even if they do not always use explicit causal language. Descriptive papers center their contribution on measurement or characterization, for example by introducing new indicators, establishing baseline levels, or documenting group differences, without developing or testing a causal mechanism. Predictive papers focus on forecasting an outcome or identifying variables that best predict it, typically using out-of-sample or cross-validated performance as an evaluation criterion, and do not interpret those variables as causes. In cases where papers combine elements of these goals, we code the general goal based on the primary contribution. Because explicit causal designs are primarily used in explanatory work, we restrict our coding of primary research design to papers classified as explanatory.

### v2

@@ Abstract
How has the credibility revolution shaped political science? We address this question by classifying 91,632 articles published between 2003 and 2023 across 156 political science journals using large language models, focusing on research design, credibility-enhancing practices, and citation patterns. We find that design-based studies—those leveraging plausibly exogenous variation to justify causal claims—have become increasingly common and receive a citation premium. In contrast, model-based approaches that rely on strong modeling assumptions have declined. Yet the rise of design-based work is uneven: it is concentrated in top journals and among authors at highly ranked institutions, and it is driven primarily by the growth of survey experiments. Other credibility-enhancing practices that help reduce false positives and false negatives, such as placebo tests and power calculations, remain rare. Taken together, our findings point to substantial but selective change—more consistent with a partial reform than a revolution.
@@ Introduction
Over the last two decades, political science has been influenced by what many describe as a credibility revolution: an intellectual movement advocating a change in how researchers conceptualize and conduct causal empirical scholarship (Angrist and Pischke 2010; Dunning 2012; Gerber and Green 2012; Morgan and Winship 2014; Samii 2016; Aronow and Miller 2019; Blair et al. 2023; Slough and Tyson 2024; Imbens and Xu 2025). But how far has this revolution actually reshaped the field? In this paper, we examine whether and to what extent the credibility revolution movement has altered the landscape of empirical causal research in political science.
@@
As an epistemic movement that reshapes what researchers view as credible evidence when assessing causal claims, the credibility revolution carries transformative potential. Its intellectual core is a reconceptualization of causality. This reconceptualization encourages a shift away from model-based strategies—which often rely on strong functional form assumptions and implicit views of treatment assignment—toward design-based approaches that exploit features of treatment assignment and rest on explicit identification assumptions. Notably, this movement views credibility not solely as a consequence of adopting specific designs such as experiments or natural experiments. In addition, it proposes a clearer articulation of identification assumptions and supporting evidence through credibility-enhancing practices such as pre-analysis plans, power analyses, and placebo tests (Sekhon 2009).
@@
Whether, on balance, this movement has produced an epistemic revolution in political science depends on how widely and deeply these ideas have reshaped actual research practice. Quantifying the depth and breadth of the movement’s influence presents two measurement challenges. The first is scope. What constitutes a sufficiently representative sample of political science to allow statements about the discipline as a whole? Previous research examining methodological changes in other social science disciplines has focused on elite journals or publishing outlets (e.g., Currie et al. 2020; Goldsmith-Pinkham 2024). However, the discipline extends far beyond these outlets, which may publish systematically different research than the rest of the field. Consequently, a systematic assessment of potential changes requires analyzing papers across multiple journal tiers over a substantial period. We address this by analyzing 91,632 articles from 156 political science journals spanning 2003–2023 using large language models (LLMs). This represents, to our knowledge, the most comprehensive analysis of political science research assembled to date.
@@
The second challenge is operationalization. The credibility revolution is an intellectual movement that seeks to transform empirical research. Translating such ideas into concrete research decisions is not straightforward. What changes should we expect to observe if its ideas have taken hold? We focus on three dimensions where change should be evident: (1) the methodological approaches employed in quantitative political science papers, (2) the adoption of practices that enhance credibility beyond research design, and (3) the professional rewards these approaches receive relative to the approaches they aim to substitute.
@@
We leverage advances in LLMs to classify our corpus along these dimensions. We begin by mapping papers by methodology and research goals, then narrow our focus to explanatory empirical quantitative papers—the subset most likely to be shaped by the credibility revolution. Within this subset, we distinguish between design-based approaches, which most straightforwardly embody the principles of the credibility revolution by leveraging features of the data-generating process for causal identification, and model-based approaches, which ground identification primarily in parametric or other modeling assumptions. We also code the use of several ancillary practices— e.g., power analysis, and placebo tests—which have the potential to further buttress the credibility of causal claims made in quantitative papers.
@@
We first assess the extent to which design-based empirical research has substituted model-based approaches in quantitative studies that aim to explain social phenomena. We find substantial, though uneven, substitution. Design-based methods increased from 15% to 40% of explanatory quantitative studies between 2003 and 2023, while model-based approaches declined from 57% to 39%. Notably, survey experiments, which narrowly identify the causal effect of information provision, accounted for 44% of all design-based papers in 2023. Further, such replacement is primarily concentrated in top journals. Taken together, these patterns suggest that while a methodological shift has occurred, it remains partial and uneven: traditional strategies continue to play a central role in explanatory research, especially in lower-visibility outlets.
@@
Second, we examine the adoption of practices that enhance credibility beyond research design. Such practices remain uncommon even among design-based studies. Pre-analysis plans and power analyses appear in only 6% and 16% of experimental papers, respectively, and placebo tests are reported in just 22% of design-based observational studies. These patterns indicate that the credibility revolution’s influence beyond methodological substitution remains limited and has not yet produced broader shifts in research standards.
@@
Third, we assess whether design-based research receives different professional rewards than model-based approaches. Descriptively, we find that design-based studies receive more citations than model-based papers, and this gap has widened over time. This citation premium has grown but now appears to have stabilized, suggesting that the discipline’s reward structure has shifted to favor more credible methods and may be reaching a new equilibrium.
@@
Overall, we find evidence of a moderate influence of the credibility revolution. Design-based methods have grown while model-based approaches have declined, indicating an increase in quantitative research that aligns with the movement’s principles. Yet the movement’s influence is neither wide nor deep. It is not wide for two reasons. First, much of political science lies outside the domain where the credibility revolution’s lessons most directly apply: quantitative studies that seek to identify causal relationships. Thus, limits to the movement’s reach partly reflect enduring methodological diversity in the discipline. Second, traditional model-based approaches remain deeply embedded in quantitative empirical explanatory research.
The movement’s reach is not deep in the sense that, even among scholars adopting design-based methods, broader credibility-enhancing practices remain rare, and the shift is driven largely by one approach—survey experiments, which accounted for nearly half of all design-based papers in 2023. In this light, the credibility revolution is best understood as a reform: a real but limited transformation, concentrated in relatively small—though not unimportant—segments of the discipline.
One might be tempted to infer from our results that political science research has become more credible because of this reform process. Design-based approaches have indeed become more prevalent, consistent with what proponents of the credibility revolution advocated. However, our analysis is descriptive, not prescriptive. It documents changes in how scholars conceptualize causality, how they justify causal claims, and what ancillary evidence they report in support of those claims, taking papers’ reporting at face value. We neither can nor intend to adjudicate whether identifying assumptions in any given paper are warranted, whether analyses were executed well, or whether specific findings are correct. There is no inherent hierarchy between the two approaches; credibility derives from the validity of assumptions and the quality of implementation. In any application, both design-based and model-based approaches may rest on defensible assumptions and be implemented well, or may be misspecified, poorly executed, and unconvincing. We therefore interpret the patterns we document as a shift in the conventions by which political scientists justify identification and present causal claims in print, not as evidence that published findings are now more likely to be correct.
@@
Our paper builds on Currie et al. (2020) and Goldsmith-Pinkham (2024), who quantify the credibility revolution movement in economics using papers published in elite journals and National Bureau of Economic Research (NBER) working papers series. A key limitation of these studies is that NBER affiliates do not represent the economics discipline as a whole. By contrast, we analyze articles from all political science journals with an impact factor above one, allowing us to examine the field well beyond the outlets targeted by scholars from elite institutions.
We also extend work by Garg and Fetzer (2025), who use LLMs to identify causal claims in economics, by systematically classifying design choices and transparency practices across an entire discipline, and complement recent large-scale diagnostic analyses of quantitative political science that evaluate the inferential properties of published research (Arel-Bundock et al. 2026). Finally, our work contributes to a broader literature on how the credibility revolution has shaped political science, including patterns of knowledge production, graduate training and hiring, and the evolution of methodological norms (Grossman et al. 2026).
The Credibility Revolution
We use the term “credibility revolution” to refer to a broad movement in the social sciences that sought to change how empirical causal research is conducted and evaluated. As an epistemic project, it foregrounded the need for explicit and defensible identification assumptions in quantitative causal research and placed particular emphasis on theorizing about the data-generating processes when justifying causal comparisons (Morgan and Winship 2014; Imbens and Rubin 2015). It emphasized the value of making identification strategies transparent and open to scrutiny, and of reducing reliance on modeling assumptions as the basis for causal claims when these are difficult to defend (Aronow and Miller 2019; Samii 2016).
Like many epistemic movements, the credibility revolution is broad, heterogeneous, and internally contested. Our aim is not to adjudicate its “exact” meaning, but to identify a core distinction that allows us to measure how empirical causal practice has changed in response to it. To that end, we adopt an analytically useful taxonomy that distinguishes between two broad approaches to empirical causal research. The taxonomy is imperfect — methods associated with a design-based approach may combine elements of the model-based approach (e.g., difference-in-differences)— but it captures a central difference in where causal identification is anchored: whether identification rests primarily on features of treatment assignment or on assumptions embedded in statistical models.
  • Design-based approaches ground causal identification in features of the data-generating process that structure treatment assignment in a plausibly exogenous fashion, such as randomization, institutional rules, or policy thresholds. While estimation may rely on statistical models, the validity of causal claims is understood to rest primarily on assumptions about treatment assignment and counterfactual comparisons, rather than on functional-form or distributional assumptions. Natural experiments, randomized controlled trials, instrumental variables, regression discontinuity designs, and difference-in-differences designs are canonical examples. [^1] In essence, the parallel trends assumption underlying difference-in-differences designs is a mean-independence assumption between the treatment and demeaned or first-differenced non-treated potential outcomes.
  • Model-based approaches, conversely, ground causal identification primarily on the correctness of parametric or statistical modeling assumptions—such as linearity, additivity, conditional independence, and distributional assumptions. In these approaches, causal interpretation hinges on the plausibility of modeling assumptions, which are often tied to theory or data structure, and may be plausible depending on the context. These assumptions are typically not anchored in an explicit claim about the treatment assignment process.
@@
The credibility revolution began in applied microeconomics in the early 1990s, driven by studies that sought to exploit institutional rules, spatial policy changes, and quasi-random variation to identify causal effects (e.g., Angrist and Krueger 1991; Card and Krueger 1994). It built on earlier critiques that regression-based causal claims were often fragile to specification choices and sensitive to researcher degrees of freedom (Leamer 1983), as well as program-evaluation evidence showing that conventional observational adjustments could diverge sharply from experimental benchmarks (LaLonde 1986). These contributions challenged prevailing practices that relied heavily on parametric regression models for causal interpretation, with causal claims resting largely on modeling choices. Researchers typically modeled the conditional expectation of an outcome using treatment variables and covariates (often without a clear distinction between the two) and then interpreted regression coefficients as causal effects. This practice placed a heavy burden on modeling assumptions, often without direct scrutiny of the underlying identification strategy.
These novel empirical developments built on formal frameworks for causal inference in statistics that emphasized explicit statements of identification assumptions. The Neyman–Rubin potential outcomes framework provided a common language for defining causal effects (Rubin 1974; Holland 1986), while the structural causal model and directed acyclic graph framework provided a graphical representation of causal assumptions (Pearl 2009). Additionally, related contributions formalized the assumptions underlying instrumental-variable designs and local average treatment effects (Imbens and Angrist 1994).
Several research designs became especially prominent within the movement. Difference-in-differences designs provided a flexible framework for exploiting policy changes and temporal variation (Card and Krueger 1994). Natural experiments leveraged institutional rules or geographic discontinuities to isolate plausibly exogenous variation in treatment assignment (Dell 2010), while instrumental variable approaches addressed unobserved confounding under explicit exclusion assumptions (Acemoglu et al. 2001). Randomized controlled trials gained renewed attention for their transparency in identifying average treatment effects, and regression discontinuity designs offered credible identification in narrow windows around policy and electoral thresholds (Lee and Lemieux 2010). Despite differences in scope and assumptions, these designs share an emphasis on explicitly linking causal identification to features of the empirical context.
Political science absorbed these ideas, with some lag, through multiple channels. Early field experiments demonstrated the feasibility of randomized interventions in political settings (Gerber and Green 2000; Wantchekon 2003), while the synthesis of natural experiments in Dunning (2012) helped codify design-based reasoning within the discipline. Regression discontinuity designs became influential in electoral studies (Eggers et al. 2015), and instrumental-variable approaches grounded in local average treatment effects gained traction in comparative and international research (Sovey and Green 2011). Difference-in-differences designs also became increasingly common, particularly in studies of policy change (Bechtel and Hainmueller 2011).
The diffusion of credibility-revolution ideas also coincided with increased attention to research practices intended to make empirical claims more transparent and, thereby, to increase the credibility of reported findings. Robustness checks, placebo tests, and sensitivity analyses became more common, alongside the spread of pre-analysis plans and data and code sharing. While transparency and reproducibility practices have distinct intellectual origins from design-based identification strategies, they similarly emphasize making assumptions and analytic choices visible, ensuring their soundness, and thereby making them more credible.
These developments have not gone uncontested. A growing body of scholarship has emphasized the limitations and trade-offs of design-based approaches. One critique concerns the locality of estimands: credible causal estimates often pertain to narrow subpopulations or institutional settings, limiting their theoretical and policy relevance (e.g., Heckman and Smith 1995). A related concern is that the focus on quasi-random variation can narrow the scope of inquiry. In privileging what can be credibly identified, researchers may sideline substantively central but historically contingent processes—such as democratization or state formation—that resist design-based identification (Mahoney and Goertz 2006).
Other critiques highlight the tension between credible identification and interpretation. Even when causal effects are well identified, their meaning often rests on untested assumptions about mechanisms, scope conditions, and counterfactual relevance. These concerns are especially salient when generalizing from localized interventions to broader theories of political behavior or institutional change (Findley et al. 2021). Taken together, these critiques underscore that the credibility revolution reflects a specific epistemological orientation—one that prioritizes internal validity and research design rigor—but does not resolve longstanding tensions among identification, explanation, and generalization in empirical political science.
On Survey Experiments
Survey experiments occupy an ambiguous position in the genealogy of the credibility revolution. They emerge from a distinct epistemic lineage—rooted in psychology and behavioral science—that predates the methodological debates over causal identification that animated applied microeconomics in the 1990s and 2000s. Many early survey experiments were developed primarily as tools for measurement and elicitation, designed to probe latent attitudes, beliefs, or cognitive processes (Mutz 2011). In this tradition, the goal was not necessarily to estimate the causal effects of real-world political interventions, but rather to improve our understanding of the psychological mechanisms underlying political behavior. In this respect, survey experiments are not direct descendants of the credibility revolution as it coalesced around the identification of causal effects in field settings.
At the same time, survey experiments are experiments and, in that sense, they do resonate with the credibility revolution’s emphasis on exogenous treatment assignment and transparency in the data-generating process. Yet the scope of the intervention in survey experiments is typically narrow: what is manipulated is information—delivered via vignettes, prompts, or framing—rather than political conditions, policies, or institutional arrangements themselves. This allows researchers to identify the causal effect of informational treatments in controlled environments. But the causal estimand often pertains to short-term attitudinal shifts in response to textual stimuli, rather than to behavioral or institutional change. As such, the causal effects identified are typically well-defined but substantively constrained, and their real-world implications can be difficult to trace (Barabas and Jerit 2010).
For these reasons, we classify survey experiments as part of the broader design-based category in our analyses, consistent with their widespread treatment in the discipline as causally identified research designs. At the same time, we recognize that their distinct origins and the limited manipulability of their treatments raise important questions about how they contribute to the goals of the credibility revolution. To reflect this duality, we present results both including and disaggregating survey experiments from other design-based approaches, allowing us to assess how their inclusion shapes broader inferences about methodological change in political science.
@@
If the credibility revolution reshaped empirical practice in political science, its effects should be observable in how quantitative causal research is conducted, reported, and rewarded. We assess its impact along three dimensions: methodological choices, transparency practices, and professional recognition. First, we examine shifts in research design. To the extent that the credibility revolution influenced empirical practice, we would expect a move away from approaches that rely primarily on modeling assumptions and toward those that ground causal claims in identification strategies that leverage explicit features of the data generating process. Empirically, this should manifest as growth in the use of design-based methods relative to model-based approaches broadly.
Second, we evaluate changes in transparency (e.g., PAPs) and reporting practices (e.g., power calculations and placebo tests). While, as mentioned, these practices have distinct intellectual roots, they are now frequently bundled with credibility-revolution standards in contemporary empirical work (Blair et al. 2023) and their proponents emphasize their ability to lend credibility to causal claims (Sekhon 2009). If such practices have diffused beyond research design alone, we should observe increased adoption of practices that clarify identification assumptions and analytic decisions.
Third, we assess whether these methodological shifts are reflected in professional rewards. If standards of credible causal inference have changed within the discipline, research that employs design-based methods or associated transparency practices may receive greater recognition. We examine this by analyzing publication patterns and citation outcomes, testing whether design-based studies are systematically rewarded relative to model-based work, holding constant publication timing and venue.
@@
We evaluate these predictions using a corpus of political science research from the past two decades. In this section, we describe the data, key measurements, and methods used to construct these measures.
We began with a set of 188 journals identified by Clarivate as part of the political science discipline and having a Scientific Journal Rankings (SJR) score of at least one. This threshold roughly corresponds to the upper tier of the field in citation influence and focuses attention on outlets that sit near the center of political science’s citation network. From this list, we removed journals for three distinct reasons: indexing (three journals not indexed in Scopus), quality control (three journals that lacked peer review), and language or format (seven journals that were not published in English and one listed as a book series). The resulting sample included 174 English-language, peer-reviewed, Scopus-indexed political science journals. Using Scopus, we extracted journal-level metadata, including citation statistics, and article-level information for all publications in these selected journals. This process produced a dataset of 129,751 articles published between 2003 and 2023.
@@
Next, we obtained full-text versions of the articles where possible and successfully retrieved the complete text for 91,632 articles from 156 different journals. Full-text availability varies somewhat across publishers and subfields, so the full-text corpus is broad but not exhaustive. We then used a combination of supervised machine learning and a commercial LLM, gpt-4o, to classify complete-text articles along 19 research design dimensions (see description of the approach below). Figure 1 shows the yearly number of published papers in our full-text sample, both overall and by subfield. The volume of published political science work has grown rapidly. From 1,546 papers in 2003, the count has more than quadrupled to over 8,109 in 2023.222See Grossman et al. (2026) for an analysis of the factors underlying the dramatic increase in political science publication volume.
Constructing Measures
@@
We construct a series of variables to characterize each paper’s methodological features, substantive focus, and transparency practices. Details of the coding procedure appear in Section A.2 in the Supplementary Materials. We first assign each paper a subfield label drawn from six categories: American Politics, Comparative Politics, International Relations, Methodology and Formal Theory, Political Theory and Philosophy, and Public Policy/Administration. The label reflects the paper’s primary substantive contribution.
@@
We then identify whether the paper is an empirical quantitative study. A paper is coded as empirical quantitative if it conducts its own analysis of observational or experimental data, including reanalyses of existing datasets. Simulation exercises or purely methodological discussions without data are not included. Papers whose main method is qualitative, formal-theoretic, or normative are coded as non-quantitative for our purposes. We further assign one of three general goals—descriptive, explanatory, or predictive—only to empirical quantitative papers.
@@
Explanatory papers investigate the causes or consequences of social phenomena and develop evidence consistent with one or more causal relationships, even if they do not always use explicit causal language. Descriptive papers center their contribution on measurement or characterization, for example by introducing new indicators, establishing baseline levels, or documenting group differences, without developing or testing a causal mechanism. Predictive studies focus on forecasting an outcome or identifying variables that best predict it, typically using out-of-sample or cross-validated performance as an evaluation criterion, and do not interpret those variables as causes. In cases where papers combine elements of these goals, we code the general goal based on the paper’s primary contribution. Because causal claims are generally not used in predictive and descriptive work, we restrict our coding of primary research design to papers classified as explanatory.